OpenStudy (anonymous):

Which function is even? (CHOOSE ALL THAT APPLY!!!!) A. y = cot x B. y = sin x C. y = csc x D. y = sec x **my answer: A and D is that right?? :)

OpenStudy (anonymous):

hey @Luis_Rivera :)

OpenStudy (anonymous):

hi. yes, it A and D

OpenStudy (anonymous):

awesome!! kk thanks!! :)

OpenStudy (anonymous):

kk

OpenStudy (mertsj):

(2) Odd Function A function f (t) is said to be 'Odd' when it satisfies the condition f (-t) = -f (t) For example, sin (-t) = -sin t, csc (-t) = -csc tan (-t) = -tan t, cot (-t) = -cot (t) are 'odd' circular functions.

OpenStudy (mertsj):

(1) Even Function A function f (t) is said to be 'Even' when it satisfies the condition f (-t) = f (t) For example, cos (-t) = cos t, sec (-t) = sec t are even circular functions. It should be noted that the graph of an even function is symmetrical about the vertical. (i.e. y axis.)

OpenStudy (anonymous):

wait, so does that mean A and D are still the even functions here then?

OpenStudy (anonymous):

i think there both right